OpenStudy (anonymous):

-7-3(-8x+8)=-4(-4x+2)-8

OpenStudy (anonymous):

-7-3(-8x+8)=-4(-4x+2)-8 -10(-8x+8) = 16x-8-8 80x-80=16x-16 64x=64 x=1 is this correct?

OpenStudy (anonymous):

No^^. You can't combine the -7 and the -3 AND THEN MULTIPLY. It goes against order of operations. You want to distribute the -3 then combine. It should read like this: \[-7+24x-24=16x-8-8\] \[24x-31=16x-16\] \[8x=15\] \[x=\frac{15}{8}\]

OpenStudy (anonymous):

now i get it..haha...thanks :)

OpenStudy (anonymous):

No problem xP